OpenStudy (anonymous):

I need help What is the simplified form of 1/x-1/y/1/x+1/y

OpenStudy (saifoo.khan):

is it's complete?

OpenStudy (anonymous):

what

OpenStudy (saifoo.khan):

the equation?

OpenStudy (anonymous):

yes they want you to simplify it its a complex fraction

OpenStudy (saifoo.khan):

\[\frac{\frac{1}{x}- \frac 1y}{\frac1x+\frac1y}\]It's like that right?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Taking lcm, "xy" for both, \[\Large \frac{\frac{y - x}{xy}}{\frac{x+y}{xy}}\]

OpenStudy (saifoo.khan):

\[\frac{y-x}{xy} \div \frac{x+y}{xy}\]\[\frac{y-x}{xy} \times \frac{xy}{x+y}\]

OpenStudy (saifoo.khan):

"xy" cancels out, leaving us with\[\frac{y-x}{x+y}\]

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

welcome